# Towards Generalist Robot Policies: What Matters in Building Vision-Language-Action Models

### v1

@@ Abstract
Foundation Vision Language Models (VLMs) exhibit strong capabilities in multi-modal representation learning, comprehension, and reasoning. By injecting action components into the VLMs, Vision-Language-Action Models (VLAs) can be naturally formed and also show promising performance. Existing work has demonstrated the effectiveness and generalization of VLAs in multiple scenarios and tasks. Nevertheless, the transfer from VLMs to VLAs is not trivial since existing VLAs differ in their backbones, action-prediction formulations, data distributions, and training recipes. This leads to a missing piece for a systematic understanding of the design choices of VLAs. In this work, we disclose the key factors that significantly influence the performance of VLA and focus on answering three essential design choices: which backbone to select, how to formulate the VLA architectures, and when to add cross-embodiment data. The obtained results convince us firmly to explain why we need VLA and develop a new family of VLAs, RoboVLMs, which require very few manual designs and achieve a new state-of-the-art performance in three simulation tasks and real-world experiments. Through our extensive experiments, which include over 8 VLM backbones, 4 policy architectures, and over 600 distinct designed experiments, we provide a detailed guidebook for the future design of VLAs. In addition to the study, the highly flexible RoboVLMs framework, which supports easy integrations of new VLMs and free combinations of various design choices, is made public to facilitate future research. We open-source all details, including codes, models, datasets, and toolkits, along with detailed training and evaluation recipes at: robovlms.github.io.
@@
Figure 1: This work mainly considers four questions for building VLAs based on VLMs: Why use VLMs-based VLAs; Which backbone to use; How to formulate the VLAs; and When to use cross-embodiment data as an extra data source. With our proposed RoboVLMs, we can easily transfer VLMs into generalist policies which supports multiple embodiments, various scenarios, and tasks.
@@ I Introduction
Building generalizable robot policies capable of perceiving, reasoning, and interacting with the physical environment given human instructions has been a long-standing challenge in robotics [5, 7, 4, 32]. Recently, there has been active exploration into learning robot foundation models by fine-tuning the Vision-Language Models (VLMs) on robot data with certain architectural adjustments. The resulting models, also referred to as Vision-Language-Action Models (VLAs), show promising results in both simulated and real-world tasks [7, 20, 22] [^1] Although the rigorous definition of VLAs is not consistent in different works, we regard fine-tuning pre-trained VLMs as the key factor to identify VLAs in this work.. Except for VLAs, there also exist various generalist policies, e.g., the ones from video models or even from scratch. Therefore, a natural question arises: Why do we prefer VLAs built upon large-scale pre-trained VLMs? Compared with other generalist policies, a mostly believed reason for utilizing VLM-based VLAs is that VLMs have demonstrated strong capabilities in learning generalized and robust representations of multi-modal data, such as text, images/videos, through extensive training on web-scale data. Such capabilities can inspire the adaptation of robot foundation models to bridge the gap between highly diverse open-world scenes and limited robotic data. However, it remains an open problem to what extent large-scale vision-language pre-train facilitates generalist robot policies. Moreover, a large and diverse set of different VLMs emerged rapidly with different kinds of LLM backbone, training data, model sizes, architectures, and training recipes. Which kind of VLM backbones is more suitable for robot manipulation is also a crucial issue for the development of successful VLAs.
@@
In addition to the VLA itself, the quality and diversity of the training data used to develop VLAs are equally critical. With recent progress achieved by well-known VLAs [36, 7, 20, 32, 4], large-scale data from different sources is important to further improve performance in terms of robustness and generalization against out-of-distribution tasks and environments. Yet, they differ largely in detailed training recipes: some utilize additional data to further pre-train VLMs, refining representations closer to robotic manipulation tasks [4], while others co-train VLAs alongside in-domain tasks [36, 7, 20, 32]. Moreover, by sufficiently pre-trained on diverse manipulation skills, robot policies are expected to learn new skills with minimal demonstrations [12]. Consequently, in case of developing efficient VLAs, When to leverage the large-scale cross-embodiment data becomes an intriguing issue.
@@ II Main Results and Findings
Real Robot Benchmark [8] consists of over 70K teleoperated human trajectories used to fine-tune robot policies, covering 105 manipulation tasks. To evaluate the performance of models on this benchmark, we adopt the approach outlined in [21], testing each model on one Simple setting and four challenging Unseen settings. Examples of these settings are shown in Fig.4. In total, we evaluate each VLA across 20 tasks, with 5 settings per task and 3 rollouts per setting, reporting the average success rate for each setting. A detailed description of the benchmarks is provided in Appendix J and Appendix C.
@@ II-A Why do we prefer VLAs?
Figure 5: Evaluation results on the SimplerEnv simulation benchmarks, including the WidowX+Bridge and Google Robot environments. Performance of baseline methods are referred from Li et al. [23]. The KosMos P.H. built by RoboVLMs is the best VLA structure investigated, which is trained over fixed training steps. Detailed numerical results can be further referred to Appendix G.
@@
Figure 6: Ablation study of VLAs for vision-language pre-train on different settings of the CALVIN benchmark. “P.H." denotes policy head. “No VL" suggests models without VL pre-training. “5x" represents training with 5x re-generated training data. Full results on different training setting and data scale are shown in Appendix D.
We also investigated the impact of vision-language pre-training on the generalization and data efficiently (Fig. 6 and Tab. IV), and the detailed result in shown in Appendix G. For the generalization in CALVIN, we adopt the official setting: training models on the split of ABC and validating performance on D. To evaluate data efficiency, we conduct experiments on model scales ranging from 3B to 9B and various data scales: 10% training data (0.1x ABCD) the standard setting (ABCD), and 500% training data (5x ABCD). The additional data originates from the officially released unlabeled datasets, following the setups as introduced in Wu et al. [43], the detailed results on different data scales are shown in Appendix D.
@@
The qualitative results are shown in Appendix J, including success rollouts in various settings and some representative failure cases. KosMos P.H. not only outperforms baseline models in basic setting tasks like Open Drawer, Pickup Egglant and Press the Toaster Switch, but also achieves better performance over unseen objects, distractors, and backgrounds. Furthermore, as shown in Fig. 8, KosMos P.H. emerges with self-correction ability, it can realize the incorrect positions of the end effector and correct its future trajectory to complete the task successfully. Note that this ability does not appear in the other tested baselines, and this kind of data is not contained in the training dataset.
@@ III Discussion
This empirical study mainly focuses on what matters in building Visual-Language Agents (VLAs) from a pre-trained Vision-Language Model (VLM). We first raise four essential questions for building a VLM-based VLA: Why do we need VLAs instead of other generalist policies, and by outperforming the existing methods by a large margin, we illustrate the necessity of studying VLAs. For the next step, we describe the key components for building a VLM-based VLA: Which kind of VLM backbone to utilize, How to train the model to generate action, and When should we add cross-embodiment data into training stages. To answer these questions, we built a unified framework for a fair comparison of existing VLAs and designed a series of bottom-up systematic experiments. To answer these questions, we conduct extensive experiments across three simulators and more than 240 rollouts within 20 tasks in real-world scenarios, and we can conclude from the experiments that: For the Why question, utilizing VLM-based VLAs could achieve high performance and generalization, and is a promising path to generalist robotics policy; For the Which problem, we find that VLMs with sufficient vision-language pre-training over large scale vision-language datasets is suitable for constructing VLAs. Note that for VLAs built upon Qwen-VL and LLaVA, the performance is surprisingly low, comparing with their original performance on vision-language tasks. After adding a perciever resampler after the vision encoder, we find that the VLAs based on Qwen-VL and LLaVA could obtain great performance gain and reach reasonable performance. We hypothesis that the performance gain is related to the image resolution and number of vision tokens in the input token sequence. However, Different VLMs have different kinds of training data, architecture, model size, data scale, LLM backbones, training recipes, etc. Nevertheless, the selected 8 VLM backbones are insufficient to reveal the relationships between these important components of a pre-trained VLM and VLA. It will be a valuable future direction to further investigate such relationships; For the How problem, we investigate the performance, generalization, and data efficiency of different VLA structures, and find that integrating history observations is essential for VLAs, and policy head is a more effective and efficient history aggregating method compared with interleaved; For the When problem, we compare three training recipes with cross-embodiment integrated at different stages, and conclude that extra in-domain data shows beneficial, and large-scale cross-embodiment pre-training further improves overall as well as few-shot performance. Besides answering the raised questions and finding out what matters for building VLM-based VLAs, we build RoboVLMs, an easy-to-use framework for constructing VLM-based VLAs. We will open-source the RoboVLMs codebase with detailed guidelines, the checkpoints of the RoboVLMs built VLAs, and the real-world collected dataset. We anticipate that our research will bolster the open-source community and expedite progress in the realms of vision-language learning and foundational models for robotics.
@@ IV-E Discussions about Structures
Appendix A Contributions
@@
Appendix B Implementation Details
@@
Appendix C Benchmark Details
@@
Appendix D Detailed Performance on CALVIN
@@
Appendix E Diverse Backbone
@@ Appendix F Diverse ph
Appendix G Detailed Performance on SimplerEnv
@@
RoboVLMs (Ours) 0.920 0.810 0.980 0.903 0.625 0.185 0.463 0.324 0.241
Appendix H Sub-task Performance with Cross-Embodiment Dataset
@@
Appendix I Rollout Examples in SimplerEnv
@@
Appendix J Rollout Examples in Real-World Experiements

### v2

@@ Abstract
Foundation Vision Language Models (VLMs) exhibit strong capabilities in multi-modal representation learning, comprehension, and reasoning. By injecting action components into the VLMs, Vision-Language-Action models (VLAs) can be naturally formed and also show promising performance. Existing work has demonstrated the effectiveness and generalization of VLAs in multiple scenarios and tasks. Nevertheless, the transfer from VLMs to VLAs is not trivial since existing VLAs differ in their backbones, action-prediction formulations, data distributions, and training recipes. This leads to a missing piece for a systematic understanding of the design choices of VLAs. In this work, we disclose the key factors that significantly influence the performance of VLA and focus on answering three essential design choices: which backbone to select, how to formulate the VLA architectures, and when to add cross-embodiment data. The obtained results convince us firmly to explain why we prefer VLA and develop a new family of VLAs, RoboVLMs, which require very few manual designs and achieve a new state-of-the-art performance in three simulation tasks and real-world experiments. Through our extensive experiments, which include over 8 VLM backbones, 4 policy architectures, and over 600 distinct designed experiments, we provide a detailed guidebook for the future design of VLAs. In addition to the study, the highly flexible RoboVLMs framework, which supports easy integrations of new VLMs and free combinations of various design choices, is made public to facilitate future research. We open-source all details, including codes, models, datasets, and toolkits, along with detailed training and evaluation recipes at: robovlms.github.io.
@@
Figure 1: This work mainly considers four questions for building VLAs based on VLMs: Why do we prefer VLAs; Which backbone to use; How to formulate the VLAs; and When to use cross-embodiment data as an extra data source. With our proposed RoboVLMs, we can easily transfer VLMs into generalist robot policies that support multiple embodiments, various scenarios, and tasks.
@@ I Introduction
Building generalizable robot policies capable of perceiving, reasoning, and interacting with the physical environment given human instructions has been a long-standing challenge in robotics [5, 7, 4, 35]. Recently, there has been an active exploration into learning robot foundation models by fine-tuning the Vision-Language Models (VLMs) on robot data with certain architectural adjustments. The resulting models, also referred to as Vision-Language-Action Models (VLAs), show promising results in both simulated and real-world tasks [7, 22, 24] [^1] Although the rigorous definition of VLAs is not consistent in different works, we regard fine-tuning pre-trained VLMs as the key factor to identify VLAs in this work.. Except for VLAs, there also exist various generalist policies, e.g., the ones from video models or even from scratch. Therefore, a natural question arises: Why do we prefer VLAs built upon large-scale pre-trained VLMs? Compared with other generalist policies, a mostly believed reason for utilizing VLM-based VLAs is that VLMs have demonstrated strong capabilities in learning generalized and robust representations of multi-modal data, such as text, images/videos, through extensive training on web-scale data. Such capabilities can inspire the adaptation of robot foundation models to bridge the gap between highly diverse open-world scenes and limited robotic data. However, it remains an open problem to what extent large-scale vision-language pre-train facilitates generalist robot policies. Moreover, a large and diverse set of different VLMs emerged rapidly with different kinds of LLM backbone, training data, model sizes, architectures, and training recipes. Which kind of VLM backbones is more suitable for robot manipulation is also a crucial issue for the development of successful VLAs.
@@
In addition to the VLA itself, the quality and diversity of the training data used to develop VLAs are equally critical. With recent progress achieved by well-known VLAs [39, 7, 22, 35, 4], large-scale data from different sources is important to further improve performance in terms of robustness and generalization against out-of-distribution tasks and environments. Yet, they differ largely in detailed training recipes: some utilize additional data to further pre-train VLMs, refining representations closer to robotic manipulation tasks [4], while others co-train VLAs alongside in-domain tasks [39, 7, 22, 35]. Moreover, by sufficiently pre-trained on diverse manipulation skills, robot policies are expected to learn new skills with minimal demonstrations [13]. Consequently, in the case of developing efficient VLAs, When to leverage the large-scale cross-embodiment data becomes an intriguing issue.
@@ II Main Results and Findings
Real Robot Benchmark [8] consists of over 70K teleoperated human trajectories used to fine-tune robot policies, covering 105 manipulation tasks. To evaluate the performance of models on this benchmark, we adopt the approach outlined in [23], testing each model on one Simple setting and four challenging Unseen settings. Examples of these settings are shown in Fig.4. In total, we evaluate each VLA across 20 tasks, with 5 settings per task and 3 rollouts per setting, reporting the average success rate for each setting. A detailed description of the benchmarks is provided in Appendix K and Appendix D.
@@ II-A Why do we prefer VLAs?
Figure 5: Evaluation results on the SimplerEnv simulation benchmarks, including the WidowX+Bridge and Google Robot environments. Performance of baseline methods are referred from Li et al. [25]. The KosMos P.H. built by RoboVLMs is the best VLA structure investigated, which is trained over fixed training steps. Detailed numerical results can be further referred to Appendix H.
@@
Figure 6: Ablation study of VLAs for vision-language pre-train on different settings of the CALVIN benchmark. “P.H." denotes policy head. “No VL" suggests models without VL pre-training. “5x" represents training with 5x re-generated training data. Full results on different training setting and data scale are shown in Appendix E.
We also investigated the impact of vision-language pre-training on the generalization and data efficiently (Fig. 6 and Tab. IV), and the detailed result in shown in Appendix H. For the generalization in CALVIN, we adopt the official setting: training models on the split of ABC and validating performance on D. To evaluate data efficiency, we conduct experiments on model scales ranging from 3B to 9B and various data scales: 10% training data (0.1x ABCD) the standard setting (ABCD), and 500% training data (5x ABCD). The additional data originates from the officially released unlabeled datasets, following the setups as introduced in Wu et al. [47], the detailed results on different data scales are shown in Appendix E.
@@
The qualitative results are shown in Appendix K, including success rollouts in various settings and some representative failure cases. KosMos P.H. not only outperforms baseline models in basic setting tasks like Open Drawer, Pickup Egglant and Press the Toaster Switch, but also achieves better performance over unseen objects, distractors, and backgrounds. Furthermore, as shown in Fig. 8, KosMos P.H. emerges with self-correction ability, it can realize the incorrect positions of the end effector and correct its future trajectory to complete the task successfully. Note that this ability does not appear in the other tested baselines, and this kind of data is not contained in the training dataset.
@@ III Discussion
This empirical study mainly focuses on what matters in building Visual-Language-Action models (VLAs). We raise four essential questions for building a VLA: Why do we need VLAs instead of other generalist policies, and by outperforming the existing methods by a large margin, we illustrate the necessity of studying VLAs. For the next step, we describe the key components for building a VLM-based VLA: Which kind of VLM backbone to utilize, How to train the model to generate action, and When should we add cross-embodiment data into training stages. To answer these questions, we built a unified framework for a fair comparison of VLAs and designed a series of bottom-up systematic experiments. To answer these questions, we conduct extensive experiments across three simulators and more than 240 rollouts within 20 tasks in real-world scenarios, and we can conclude from the experiments that: For the Why question, VLAs could achieve high performance and generalization, and is a promising path to generalist robotics policy; For the Which problem, we find that VLMs with sufficient vision-language pre-training over large scale vision-language datasets is suitable for constructing VLAs. For the How problem, we can investigate the performance, generalization, and data efficiency of different VLA structures, and find that integrating history observations is essential for VLAs, and policy head is a more effective and efficient history aggregating method compared with interleaved; For the When problem, we compare three training recipes with cross-embodiment integrated at different stages, and conclude that extra in-domain data shows beneficial, and large-scale cross-embodiment pre-training further improves overall as well as few-shot performance. As a byproduct of answers to the raised questions, we built an easy-to-use framework for easily integrating arbitrary VLMs and turning them into VLAs, named RoboVLMs.
Under investigated observations. During our experiments, we found that VLAs built upon Qwen-VL and LLaVA, the performance is surprisingly low, compared with their original performance on vision-language tasks. After adding a perceiver resampler after the vision encoder, we found that the VLAs based on Qwen-VL and LLaVA could obtain great performance gain and reach reasonable performance. We hypothesize that the performance gain is related to the image resolution and number of vision tokens in the input token sequence.
Limitations. Although we make every effort to investigate the key challenges in building Vision-Language Agents (VLAs), this work remains preliminary and has several limitations at the current stage. (1) For the sake of quick and simple expansion over existing Vision-Language Models (VLMs), we retain the multi-modal interaction structure within the VLM (e.g., attention masks, mixture of experts). On top of this, we further develop the interaction between vision, language, and actions, which is a common approach in most existing works [22, 24]. However, a specialized design for the architecture and multi-modal interaction with actions has the potential to yield superior performance (e.g., the $\pi_{0}$ model [4]), and warrants further exploration. (2) The categorizations and formulations of VLAs considered here are simplified and limited for the reasons outlined in (1). (3) The action tokenization, policy head, and corresponding training objectives are not fully explored in this work. For example, techniques like VQ-VAE [42], diffusion models [17, 9], and flow matching [26, 12, 4] remain under-explored in the context of VLAs. (4) The set of VLM backbones considered in this study is limited and can be actively expanded. (5) Deploying such large models for real-time robotic control remains a significant challenge.
Future works. For future work, we envision several potential directions for advancing generalist robot policies. 1) As aforementioned, our current approach faces limitations in the design of internal structures for VLMs, policy heads, and corresponding training objectives. Further investigation into more fine-grained design choices for VLAs could be highly valuable, as recent studies suggest they play a significant role in improving efficiency and effectiveness [4]. 2) Beyond semantic generalization, an ideal generalist robot policy should be capable of handling long-horizon, complex task instructions (e.g., make breakfast), reasoning through executable actions step by step, and generating meaningful physical interactions with its environment (e.g., [52]). In our future work, we aim to explore the key elements required to develop policies with these advanced capabilities.
We have open-sourced our codebase with detailed guidelines, model weights of the strongest VLAs built by RoboVLMs, along with the real-world dataset used in our experiments. We anticipate that our research will bolster the community and expedite progress in the realms of vision-language learning and foundational models for robotics.
@@ IV-E Discussions about Structures
Appendix A Acknowledgments
We thank all the members of the robotics research team at ByteDance Research for their assistance in real-world data collection, setup design, robot maintenance, and experiments. The author Minghuan Liu is supported by the ByteDance Scholarship. We also want to thank @YouJiacheng for his active and instructive discussion on X.
Appendix B Contributions
@@
Appendix C Implementation Details
@@
Appendix D Benchmark Details
@@
Appendix E Detailed Performance on CALVIN
@@
Appendix F Diverse Backbone
@@ Appendix G Diverse ph
Appendix H Detailed Performance on SimplerEnv
@@
RoboVLMs (Ours) 0.940 0.470 0.910 0.773 0.617 0.333 0.531 0.435 0.241
Appendix I Sub-task Performance with Cross-Embodiment Dataset
@@
Appendix J Rollout Examples in SimplerEnv
@@
Appendix K Rollout Examples in Real-World Experiements